OpenStudy (anonymous):

PLEASE HELP A pet shop had 8 identical fish tanks in the shape of rectangular prisms. The width and height of each tank was (x - 7) inches and the length was 8 inches more than its width. The expression below shows the maximum amount of water which all the tanks together can hold: 8(x + 1)(x - 7)2 What does the factor 8(x + 1) represent? The total height of all the tanks The total length of all the tanks The total volume of all the tanks The total base area of all the tanks

OpenStudy (anonymous):

Someone Help

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

x-7 is width so x-7 +8 = x+1 is the length of 1 tank

hartnn (hartnn):

then 8(x+1) will be length of all 8 tanks :)

OpenStudy (anonymous):

MAY YOU explain

OpenStudy (anonymous):

I am a little confused

hartnn (hartnn):

what is the width ?

OpenStudy (anonymous):

x - 7

hartnn (hartnn):

"the length was 8 inches more than its width."

hartnn (hartnn):

length = x-7+8 good so far ?

OpenStudy (anonymous):

yes

hartnn (hartnn):

x-7+8 is just x+1 isn't it ?

OpenStudy (anonymous):

yes?

OpenStudy (anonymous):

I am asking what the answer is choice a, b, c or d

OpenStudy (anonymous):

well i mean that is cheating so I want you to explain me it

hartnn (hartnn):

so x+1 is length of 1 tank and length of 8 tanks will be 8(x+1) right ? 8(x+1) is the length of all tanks

hartnn (hartnn):

which is "total length of all tanks"

OpenStudy (anonymous):

oh ok that makes sense now

OpenStudy (anonymous):

may you help me with another one

OpenStudy (anonymous):

well actually a few more

hartnn (hartnn):

ask each question in a new post so that if i am unavailable, others can help :)

OpenStudy (anonymous):

ok thank you

hartnn (hartnn):

welcome ^_^